OpenStudy (anonymous):

ay=-9.8m/s^2 v1y=0m/s ∆dy=-91.2cm (-0.912m)±1.5mm (0.0015m) ∆t=? ∆dy= v1y∆t + 1/2ay∆t^2 -0.912m±0.0015m=0∆t + 1/2(-9.8m/s2)∆t^2 -0.912m±0.0015m=-4.9m/s^2∆t^2 0.1861224499s2±0.16%=∆t2 √0.186122449s2±0.16%=√∆t^2 0.43141911s±0.4%=∆t 0.43s±0s=∆t is this correct?

OpenStudy (anonymous):

let me see V=da/dy=-9.8 m/ss*t =vly=0 y(t)=dV/dt=-1/2*9.8m/ss t^2=-91.2cm then t^2=-91.2cm*2/-9.8m=>0.186122 t=0.431419 then check -.5*9.8*t^2=0.912 m works

OpenStudy (anonymous):

okay thanks! but are my uncertainty calculations correct though? Im not sure if I'm suppose to square root the 0.16% too

OpenStudy (anonymous):

Those can be tricky, you were probably given an equation to use for them that is less complex then what you would use in research.

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/Propagation_of_uncertainty link to the equation for error prorogation.. is this what you are working with?

OpenStudy (anonymous):

Like i know when you add or subtract data you would add absolout uncertanties, and when you multiply or divide data you would add the relative percentage uncertainties, so I'm so quite sure what you do when you need to square root an uncertainty

OpenStudy (anonymous):

no I'm not working with error prorogation

OpenStudy (anonymous):

it not normally that easy, if that is what your teacher wants, they that is easy. else is is probably something like new error=1/2(olderror/number)

OpenStudy (anonymous):

If you have not done anything like that then you are probably doing something different. These equations come form standard deviation.

OpenStudy (anonymous):

Most of the time a computer will do them for you.

OpenStudy (anonymous):

okay thanks anyways